# Unitary Quadratic Quantum Gravity in 4D

### 21 Apr 2026 v1

@@ II The dual-IHO: Berry–Keating and phase space structure
where $m_{0}^{2}>0$ is mass square of the scalaron, and $m_{2}^{2}>0$ (or $\mu_{2}^{2}>0$) is the mass scale associated with the additional spin–2 mode of the theory. The theory has a massless spin-2 and a massive scalar (for $\alpha>0$) (i.e., scalaron) as on-shell (also off-shell) propagating degree of freedom. The status of additional spin-2 depends on the coefficient of the Weyl square. If $\beta<0$, the spin-2 becomes a pathological ghost and can be an on-shell mode whereas the $\beta>0$ case, it becomes a dual IHO spin-2 field, which is a purely off-shell mode as we explain below. The four systems relevant to quadratic gravity are distinguished by their Hamiltonians:
@@
The IHO and dual-IHO Hamiltonians are indefinite, not negative definite: this single fact separates them categorically from the ghost. In the canonically rotated coordinates $Q=(\tilde{p}+\tilde{q})/\sqrt{2}$, $P=(\tilde{p}-\tilde{q})/\sqrt{2}$, both take the Berry–Keating (BK) form [4]:

### 23 Apr 2026 v2

@@ II The dual-IHO: Berry–Keating and phase space structure
where $m^{2}>0$ is mass square of the scalaron, and $m_{2}^{2}>0$ (or $\mu_{2}^{2}>0$) is the mass scale associated with the additional spin-2 mode of the theory. The theory has a massless spin-2 and a massive scalar (for $\alpha>0$) (i.e., scalaron) as an on-shell (also off-shell) propagating degree of freedom. The status of additional spin-2 depends on the coefficient of the Weyl square. If $\beta<0$, the spin-2 becomes a pathological ghost and can be an on-shell mode, whereas in the $\beta>0$ case, it becomes a dual IHO spin-2 field, which is a purely off-shell mode as we explain below. The four systems relevant to quadratic gravity are distinguished by their Hamiltonians:
@@
The IHO and dual-IHO Hamiltonians are indefinite, not negative definite. This single fact separates them categorically from the ghost. In the canonically rotated coordinates $Q=(\tilde{p}+\tilde{q})/\sqrt{2}$, $P=(\tilde{p}-\tilde{q})/\sqrt{2}$, both take the Berry–Keating (BK) form [4]: